OpenStudy (jamanorthi):

An experiment consist of rolling two fair number cubes. Find each outcomes. P(rolling a total less then 2)????

OpenStudy (anonymous):

P(sum < 2)=0

OpenStudy (jamanorthi):

How do I do that

OpenStudy (jamanorthi):

Are u there???

OpenStudy (anonymous):

sample space or exhaustive events are 11,12,13,14,15,16, 21,22,23,24,25,26, ............................, ............................., ............................., 61,62,63,64,65,66. Total no. of exhaustive events=6*3=36 smallest sum=1+1=2 largest sum=6+6=12 we do not get sum <0 no. of favourable events=0 \[Required~ probability=\frac{ 0 }{ 36 }=0\]

OpenStudy (anonymous):

correction 6*6=36

OpenStudy (jamanorthi):

Ok what if it says p(rolling a total of 9) how do I do this Is it like 9/36=1/4

OpenStudy (anonymous):

favourable events=36,45,54,63 no. of favourable events=4 P=4/36=1/9

OpenStudy (jamanorthi):

Why 4 doe?

OpenStudy (jamanorthi):

I am so confused

OpenStudy (anonymous):

you want sum=9 they are only four as i have shown above in the table.

OpenStudy (anonymous):

we have to find the numbers of favourable events i.e.,sum =9

OpenStudy (anonymous):

you got it .

OpenStudy (jamanorthi):

Yeah thanks

OpenStudy (jamanorthi):

But How did u got 36.45.54.63 And why only those number And how u got that